OpenStudy (anonymous):

A druggist has two mixtures of a certain chemical and water, one containing 10% of the chemical and the other containing 40% of the chemical. How much of each mixture should he use to make 10 ounces that contain 25% of the chemical? If x is the amount of 10% mixture used, then which expression represents the ounces of chemical contained in it?

OpenStudy (dumbcow):

x+y = 10 .1x +.4y = .25*10

OpenStudy (anonymous):

Let the solution with 25% of the chemical be solution A and the other be solution B. As given,x is 'the amount of solution A used'. Let y be the 'amount of solution B used'. Using the given data, x+y=10 and also, 25=(10*x+40*y)/(x+y) Solving both equations,we get x=5 and y=5. as for the required equation, ounces of chemical contained in final solution= (40-3x)/10